three good friends are in the same algebra class their scores on a recent test are three consecutive even integers whose sum is 216.find each score
Found 2 solutions by Alan3354, josh_jordan:
Answer by Alan3354(69443)

three good friends are in the same algebra class their scores on a recent test are three consecutive even integers whose sum is 216.find each score
----------
216/3 = 72, the middle score

Answer by josh_jordan(263)

To find each test score, let's turn this word problem into an equation. We know that all of the test scores are even numbers. We also know that when we add 2 to an even number, we get the next even number, and when we add 4 to the original even number, we get the next consecutive even number. In other words,

x + x + 2 + x + 4, where x represents the first unknown even number. We will set this equal to 216, since we know that the sum is 216:

x + x + 2 + x + 4 = 216

Then, we will combine the x's together and the numbers together on the left side of our equal sign, giving us

3x + 6 = 216

Now, subtract 6 from both sides of the equal sign. This will give us

3x = 210

Divide both sides by 3, which will give us our first test score:

x = 70

Finally, to find the other test scores, add 2 to 70, which will give us our second test score: 72. Then, add 4 to 70, which will give us our final test score: 74

Test Scores: 70, 72, 74

Validate by adding all three numbers together. 70 + 72 + 74 = 216